# Activation Function Design Sustains Plasticity in Continual Learning

### v2

@@ 1 Introduction
We argue that a more fundamental knob is hiding in plain sight: the activation function. Differences among activations often shrink in i.i.d. training once model size and optimization are tuned, but under continual, non-stationary data they can matter substantially. This motivates a property-level study of how activation shape, especially negative-side responsiveness and saturation, affects plasticity [^1] Code available at: https://anonymous.4open.science/r/activations_plasticity-E431/. Previous work has shown that alternative activations such as CReLU Shang et al. (2016); Abbas et al. (2023); Lee et al. (2023) or rational functions Molina et al. (2019); Delfosse et al. (2021b; a) can improve performance under non-stationary conditions. Section 2 closes with a simple IID vs. class-incremental (C-IL) comparison using a shared setup to illustrate this contrast and motivate the case studies that follow.
@@ 5.1 Smooth-Leaky and Randomized Smooth-Leaky
This randomized variant preserves the strict floor and $C^{1}$ transition while encouraging robustness to small variations in negative-side responsiveness. Limitations of multi-parameter design are explained in App. B.2.
@@ Acknowledgments
This material is based on work supported by the National Science Foundation under Grant No. 2218063. The authors acknowledge the Vermont Advanced Computing Center (VACC) at the University of Vermont for providing computational resources that have contributed to the research results reported in this paper.
@@ B.4 Hyper-parameter Sweeps for Activation Functions
ReLU / CReLU $\alpha$ 0.0

### v3

# arXiv:2509.22562v3[cs.LG] 26 Feb 2026

@@ 1 Introduction
We argue that a more fundamental knob is hiding in plain sight: the activation function. Differences among activations often shrink in i.i.d. training once model size and optimization are tuned, but under continual, non-stationary data they can matter substantially. This motivates a property-level study of how activation shape, especially negative-side responsiveness and saturation, affects plasticity [^1] Code available at: https://github.com/lute47lillo/activations_plasticity. Previous work has shown that alternative activations such as CReLU Shang et al. (2016); Abbas et al. (2023); Lee et al. (2023) or rational functions Molina et al. (2019); Delfosse et al. (2021b; a) can improve performance under non-stationary conditions. Section 2 closes with a simple IID vs. class-incremental (C-IL) comparison using a shared setup to illustrate this contrast and motivate the case studies that follow.
@@ 5.1 Smooth-Leaky and Randomized Smooth-Leaky
This randomized variant preserves the strict floor and $C^{1}$ transition while encouraging robustness to small variations in negative-side responsiveness. Limitations of multi-parameter design and computational budget-fairness are explained in App. B.2.
@@ Acknowledgments
This material is based on work supported by the National Science Foundation under Grant No. 2218063 and 2239691. The authors acknowledge the Vermont Advanced Computing Center (VACC) at the University of Vermont for providing computational resources that have contributed to the research results reported in this paper.
@@ Appendix B Experimental and Architectural Design
Budget-matched robustness and computational fairness.
Our proposed activation family admits a larger hyperparameter grid than some baselines (because it has more tunable parameters), potentially inflating performance by enabling a broader search. To address this, we evaluate budget-matched robustness rather than enforcing an arbitrary equality in the raw number of tested settings across methods. For each activation $i$ and dataset, we define a tuning budget $N_{i}$ as the number of evaluated hyperparameter configurations for $i$ (Table B2). We then simulate a practitioner with budget $N_{i}$ who can try $N_{i}$ randomly chosen configurations and keep the best result: we bootstrap a best-of-$N_{i}$ distribution by repeatedly sampling $N_{i}$ configurations (with replacement) and taking the maximum accuracy. We perform a one-sided Mann–Whitney U test with alternative
$H_{1}:\ \mathrm{BestOf}_{N_{i}}(\text{Rand Smooth-Leaky})>\mathrm{BestOf}_{N_{i}}(i),$
and report the resulting p-values in Table B1. Under this formulation, p-values close to 0 (reported as $10^{-4}$ if values are smaller than such for simplicity) indicate that Rand. Smooth-Leaky yields consistently higher best-of-$N_{i}$ outcomes than activation $i$ when both are given the same number of random hyperparameter trials, while p-values closer to $1$ indicate the opposite direction.
The results show that Rand Smooth-Leaky remains competitive under matched budgets across a wide range of baselines and datasets. In particular, for Permuted-MNIST, Random Label MNIST, Random Label CIFAR, and Continual ImageNet, Rand Smooth-Leaky significantly dominates most baselines under their own budgets (e.g., ReLU, Sigmoid, Tanh, PReLU, Swish/SiLU, GeLU, eLU, SeLU, CReLU, Rational, and SwiGLU yield $p<10^{-4}$ in nearly all cases). This indicates that the strong performance of Rand Smooth-Leaky is not contingent on conducting an unusually large, constrained hyperparameter search: even when restricted to the same tuning budget as each baseline (typically $N_{i}\in\{2,36,40,62,64\}$ depending on the activation), random budgeted sampling is sufficient to recover high-performing configurations with high probability.
Table B1 also makes the failure modes explicit via the directionality of the one-sided test. DeepFourier yields p-values of $1.00$ or close to $1.00$ on several datasets (e.g., Permuted-MNIST, and CIFAR 5+1), indicating that under its small budget ($N_{i}=2$) its best-of-$N_{i}$ can exceed Rand Smooth-Leaky’s best-of-$N_{i}$ in those settings; conversely, Rand Smooth-Leaky significantly exceeds DeepFourier on Random Label MNIST, Random Label CIFAR, and Continual ImageNet ($p<10^{-4}$). Similarly, Smooth-Leaky exhibits a mixed pattern: Rand Smooth-Leaky is not higher on Permuted-MNIST, Random Label CIFAR, and Continual ImageNet (p-values $1.00$), but it is significantly higher on Random Label MNIST and CIFAR 5+1 ($p<10^{-4}$). However, this suggests the strength of the other activation function introduced: Smooth-Leaky. Thus, making such comparison between them independently good to whichever is better. Finally, CeLU shows a single notable exception on CIFAR 5+1 (p-value $1.00$) while being dominated elsewhere, illustrating that dataset-specific interactions can invert the relative best-of-budget ordering. While other activation functions are close but not statistically significant in some datasets.
Overall, these budget-matched tests for computational fairness directly answer the concern: our conclusions do not rely on matching the raw number of hyperparameter combinations across methods (which is not meaningful once baselines plateau within their own hyperparameter ranges). Instead, by evaluating best-of-$N_{i}$ under matched budgets derived from each baseline’s sweep size, we show that Rand Smooth-Leaky is typically robust to hyperparameter choice and achieves high performance even under limited random search budgets, while transparently identifying the few dataset/activation pairs where this dominance does not hold.
Activations Permuted MNIST Random Label MNIST Random Label CIFAR CIFAR 5+1 Continual ImageNet
ReLU $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
Leaky-ReLU $1.00$ $0.0005$ $0.5820$ $0.0004$ $10^{-4}$
Sigmoid $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
Tanh $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
RReLU $0.0310$ $0.471$ $0.0118$ $1.00$ $10^{-4}$
PReLU $10^{-4}$ $0.071$ $10^{-4} <$$10^{-4} <$$10^{-4}$
Swish (SiLU) $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
GeLU $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
CeLU $10^{-4} <$$10^{-4} <$$^{-4} 1.001$.00 $10^{-4}$
eLU $10^{-4} <$$10^{-4} <$$10^{-4}$ $0.62$ $10^{-4}$
SeLU $10^{-4} <$$10^{-4} <$$10^{-4}$ $0.37$ $10^{-4}$
CReLU $0.0008$ $0.035$ $10^{-4} <$$10^{-4} <$$10^{-4}$
Rational $0.28$ $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
SwiGLU $10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4} <$$10^{-4}$
DeepFourier $1.00$ $10^{-4}$ $0.77$ $10^{-4} <$$10^{-4}$
Smooth-Leaky $1.00$ $^{-4} 1.001$.00 $^{-4} 1.001$.00
Table B1: Budget-fairness comparison significance test. For each baseline activation $i$, combined learning rates, and dataset, let $N_{i}$ be the number of evaluated hyperparameter configurations for $i$ (Table B2). We bootstrap best-of-$N_{i}$ performance by repeatedly sampling $N_{i}$ configurations (with replacement) and taking the maximum accuracy. We then perform a one-sided Mann–Whitney U test with alternative $H_{1}:\mathrm{BestOf}_{N_{i}}(\text{Rand Smooth-Leaky})>\mathrm{BestOf}_{N_{i}}(i)$. Reported values are p-values. We perform 50 bootstrap repetitions.
Activations Permuted MNIST Random Label MNIST Random Label CIFAR CIFAR 5+1 Continual ImageNet
CeLU 64 64 64 64 64
CReLU 2 2 2 2 2
DeepFourier 2 2 2 2 2
eLU 64 64 64 64 64
GeLU 64 64 64 64 64
Leaky-ReLU 64 64 64 64 64
PReLU 36 36 36 36 36
Rational 20 26 12 41 14
ReLU 2 2 2 2 2
RReLU 40 40 40 40 40
SeLU 62 62 62 62 62
Sigmoid 2 2 2 2 2
Smooth-Leaky 175 175 175 175 175
SwiGLU 2 2 2 2 2
Swish (SiLU) 64 64 64 64 64
Tanh 2 2 2 2 2
Table B2: Configuration counts for budget fairness comparison. $N_{i}$ denotes the number of successfully evaluated hyperparameter configurations for activation $i$ on each dataset. These values define the tuning budget used in Table B1.
@@ B.4 Hyper-parameter Sweeps for Activation Functions
ReLU / CReLU / Deep Fourier $\alpha$ 0.0